### v2

# Edit Once, Update Everywhere: A Simple Framework for Cross-Lingual Knowledge Synchronization in LLMs

Abstract
Knowledge editing allows for efficient adaptation of large language models (LLMs) to new information or corrections without requiring full retraining. However, prior methods typically focus on either single-language editing or basic multilingual editing, failing to achieve true cross-linguistic knowledge synchronization. To address this, we present a simple and practical state-of-the-art (SOTA) recipe Cross-Lingual Knowledge Democracy Edit (X-KDE), designed to propagate knowledge from a dominant language to other languages effectively. Our X-KDE comprises two stages: (i) Cross-lingual Edition Instruction Tuning (XE-IT), which fine-tunes the model on a curated parallel dataset to modify in-scope knowledge while preserving unrelated information, and (ii) Target-language Preference Optimization (TL-PO), which applies advanced optimization techniques to ensure consistency across languages, fostering the transfer of updates. Additionally, we contribute a high-quality, cross-lingual dataset, specifically designed to enhance knowledge transfer across languages. Extensive experiments on the Bi-ZsRE and MzsRE benchmarks show that X-KDE significantly enhances cross-lingual performance, achieving an average improvement of +8.19%, while maintaining high accuracy in monolingual settings. The code and dataset will be available upon acceptance. Our code is available at: ahttps://github.com/YukinoshitaKaren/X_KDE.
Edit Once, Update Everywhere: A Simple Framework for Cross-Lingual Knowledge Synchronization in LLMs
Anonymous ACL submission
1 Experiments
1.1 Experimental Setup
Baselines.
Method Test in English Test in Chinese
Reliability Generality Locality Portability Reliability Generality Locality Portability Avg.
Edit in English
FT-L 53.51 50.18 94.01 53.31 51.81 51.71 85.56 55.14 61.90
FT-M 99.97 95.38 97.92 57.69 56.89 56.52 94.61 52.16 76.39
ROME 96.09 84.69 98.04 58.87 49.94 50.31 97.70 51.81 73.43
MEMIT 95.21 89.14 98.56 57.77 52.05 52.01 98.76 52.19 74.46
IKE 99.59 99.61 56.95 71.27 67.83 67.88 64.54 58.97 73.33
LTE 99.91 99.81 88.97 77.40 76.86 76.82 86.99 67.49 84.28
X-KDE(Ours) 99.93 99.87 90.15 76.41 94.81 94.65 95.05 77.43 91.04
Edit in Chinese
FT-L 40.80 40.66 94.80 55.24 54.72 53.68 66.51 48.75 56.89
FT-M 51.86 51.24 98.18 55.30 100.0 99.71 79.28 61.98 74.69
ROME 44.14 43.80 97.92 52.66 72.24 70.12 96.48 48.15 65.69
MEMIT 45.37 44.95 99.07 54.65 75.19 73.45 96.02 51.44 67.52
IKE 65.87 65.74 69.41 63.06 99.86 99.73 64.86 72.39 80.79
LTE 64.63 62.56 85.23 62.6 99.79 99.31 87.17 69.69 78.87
X-KDE(Ours) 93.49 92.22 90.56 65.55 100.00 99.11 92.85 74.14 88.49
Table 1: Cross-lingual editing performance of different methods on Llama2-chat-7B backbones. Results in green indicates the best results. “Avg.” represents the overall mean of all metrics evaluated across the two languages.
Metrics Methods en-en en-cz en-de en-du en-es en-fr en-pt en-ru en-th en-tr en-vi en-zh en-avg
Reliability FT-L 52.92 41.81 39.79 39.02 39.49 39.72 39.26 39.79 36.44 36.86 46.21 51.81 41.93
FT-M 99.96 66.93 70.16 67.17 63.69 64.98 64.22 48.96 36.46 57.54 66.80 56.89 63.65
ROME 96.36 56.54 60.82 58.89 57.41 56.43 54.91 41.69 35.44 45.76 56.94 49.94 55.93
MEMIT 95.44 62.37 64.82 64.12 59.46 61.90 58.69 44.54 36.40 49.15 61.34 52.05 59.19
IKE 99.65 83.22 80.61 79.36 76.69 78.48 75.37 67.62 54.38 76.90 81.22 67.83 76.78
LTE 100.00 84.29 81.71 80.60 77.67 79.11 77.39 72.02 62.04 78.87 81.92 76.93 79.38
X-KDE 99.93 92.78 87.43 88.89 85.71 87.49 89.87 89.32 89.66 91.23 87.55 93.07 90.24
Generality FT-L 49.60 40.75 38.87 38.36 39.68 39.12 39.56 38.97 36.89 37.18 45.89 51.71 41.38
FT-M 95.53 65.45 68.15 65.09 62.39 62.28 61.63 47.69 36.88 56.87 65.97 56.52 62.04
ROME 85.13 54.99 58.91 56.99 56.58 54.47 53.94 40.68 35.36 45.06 56.38 50.31 54.07
MEMIT 89.59 60.71 63.80 61.98 58.10 59.40 57.63 43.31 36.77 48.68 60.51 52.01 57.71
IKE 99.54 82.67 80.78 79.18 76.37 78.22 75.49 67.51 54.26 76.97 80.99 67.88 76.65
LTE 99.87 84.26 81.63 81.07 77.51 78.99 77.38 71.46 61.90 78.26 81.37 76.24 79.16
X-KDE 99.68 92.87 87.25 88.87 85.16 87.57 89.93 89.10 89.21 91.25 87.62 93.11 90.14
Table 2: Results on MzsRE dataset for editing performed in English using Llama2-7b-chat. Here, “en-zh” means that English serves as the source language and Chinese as the target language, with similar interpretations for the other pairs. “en-avg” denotes the average performance across cross-lingual scenarios.
We chose the following methods as baselines: (1) FT-L meng2022locating fine-tunes a specific layer of the feed-forward network to maximize the likelihood of target tokens; (2) FT-M zhang2024comprehensive fine-tunes the same feed-forward network layer as FT-L. Additionally, it masks the original text and applies cross-entropy loss on the target answer; (3) ROME meng2022locating employs causal mediation analysis to identify the target area for editing, and then updates the parameters of the feed-forward network layers; (4) MEMIT meng2022mass, built upon the ROME framework, enables the simultaneous update of thousands of knowledge; (5) IKE zheng2023can utilizes the in-context-learning ability of the model and provides a few-shot demonstration to guide the model’s responses based on the updated facts. (6) LTE jiang2024learning enhances the model’s instruction-following ability through supervised fine-tuning (SFT), and employs a retrieval-based mechanism to provide updated knowledge for demonstrations.
Dataset.
We evaluate X-KDE and Baselines using two widely used benchmarks: Bi-ZsRE dataset wang2023cross and MzSRE dataset wang2023retrieval. The former contains 1,037 test samples in English and an equal number in Chinese, while the latter covers twelve languages: English, Czech, German, Dutch, Spanish, French, Portuguese, Russian, Thai, Turkish, Vietnamese and Chinese, with 743 samples in each.
Backbones.
We select two public models as backbones, including LLaMA2-Chat-7B touvron2023llama and Qwen2.5-instruct-7B yang2024qwen2. These models are widely used in chatbot applications, the former excels in English, while the latter demonstrates strong multilingual abilities. For brevity, the results on Qwen2.5-instruct-7B are provided in D.2.
Figure 1: Mean batch-editing performance across four benchmarks at batch sizes 1, 10, 100, and 1000.
Figure 2: Mean sequential-editing results across four knowledge editing benchmarks, shown for data stream sizes of 1, 10, 100, 500, and 1000 (log-scale).
1.2 Results of Single Fact Editing
Table 1 and Table 2 demonstrate the main results of single fact editing, which focus on single editing cases. From these results, we can find several significant observations:
X-KDE outperforms other methods in the cross-lingual setting by a significant margin.
As shown in Table 1, when edited in English, it is evident that our method brings average performance improvements of 6.76%, compared to LTE. In particular, in the cross-lingual setting, our method achieves further performance gains in portability, which demonstrates that our method not only captures surface-level changes in wording but enables the LLM to effectively internalize the knowledge edited in the source language and apply it to the target language. In summary, LTE sets a new state-of-the-art in cross-lingual knowledge editing task.
X-KDE brings consistent and effective improvements in more complex multilingual environments.
Our method is effective not only in a bilingual Chinese-English setting but can also be generalized to additional languages. We conducted more extensive experiments on the Mzsre dataset, and the results are presented in Table 2. More detailed results in Appendix D.1. It is evident that, compared to LTE, our method exhibits 10.86% in reliability and 10.98% average gains in generality when edit in English. These results further demonstrate that, our approach significantly enhances the model’s cross-lingual abilities, enabling it to effectively apply knowledge from a pivot language to others and marking a significant step toward the democratization of knowledge.
1.3 Results of Mass Fact Editing
In the previous section, we introduced the results of single fact editing. However, real-world scenarios are often more complex, requiring simultaneous or sequential edits to multiple pieces of knowledge. Therefore, in this section, we conduct comprehensive experiments using X-KDE alongside several methods that support mass editing (FT-L, FT-M, MEMIT, and LTE) on LLaMA2-Chat-7B in both batch-edit and sequential-edit settings, and then present the corresponding results.
Test Language: en Test Language: zh
MMLU CommonSenseQA PIQA Xsum NQ CMMLU CommonSenseQA_zh CEval NQ_zh avg
LLaMA2-Chat-7B 44.78 64.21 66.43 21.24 19.39 20.64 4.67 30.08 0 30.16
 X-KDE 47.67 59.46 61.64 20.96 29.36 32.63 45.13 30.75 14.07 37.96
Table 3: General tasks performance of X-KDE, and LLaMA2-Chat-7B. Results in bold represent the best performance in each category.
X-KDE can process thousands of edits simultaneously.
In line with the single-edit procedure, we evaluate both English and Chinese edits separately. For simplicity in presentation, we take the average of these two results, as shown in Figure 1. As the batch size increases, we observe a gradual decline across all performance metrics for all method. The drop is particularly severe for MEMIT and FT-L, especially in the locality metric, which is nearly cut in half. In contrast, X-KDE achieves the best performance, maintaining the highest accuracy while exhibiting the slowest degradation rate. These results indicate that our approach remains stable in cross-lingual settings, even after thousands of edits.
X-KDE can sequential acquire new knowledge without forgetting previous information.
In the sequential-editing setting, the model integrates new knowledge on top of its previous edits, which leads to a gradual decline in performance over time. As illustrated in Figure 2, the performance of methods that modify model parameters typically degrades as the number of edits increases. For instance, MEMIT and FT-L remain stable only when the number of edits $n\leq 10$; beyond that, their performance deteriorates sharply. In contrast, knowledge storage-retrieval paradigms represented by X-KDE and LTE circumvent direct parameter modifications through external memory architectures. Moreover, X-KDE demonstrates superior cross-lingual transfer capabilities compared to LTE, achieving better performance across diverse data streams.
1.4 Results of General Tasks
A series of studies have demonstrated that knowledge editing can influence model performance across various scenarios yang2024butterfly; li2023unveiling; gu2024model. To investigate whether our method impact the model’s capabilities in unrelated domains, we conducted tests across a range of fields. Given that the cross-lingual knowledge editing task typically involves two languages, we use English and Chinese as representative examples. Multiple benchmarks are selected in these two languages, covering tasks such as commonsense reasoning, natural language understanding, open-domain QA, and general intelligence. For example, the benchmarks selected for English include MMLU hendrycks2020measuring, CommonSenseQA talmor2018commonsenseqa, PIQA bisk2020piqa, XSum narayan2018don, and Natural Questions kwiatkowski2019natural. For Chinese, the chosen benchmarks are CMMLU li2023cmmlu, CommonSenseQA_zh 2023opencompass, CEval huang2024c, and Natural Questions_zh 2023opencompass. We conducted all experiments using the OpenCompass tool 2023opencompass. The results are presented in Table 3. Overall, our method not only preserves the model’s performance in English but also significantly enhances its capabilities in Chinese. Although certain tasks, such as CommonsenseQA, XSum, and PIQA, show a decrease in performance when tested in English, the overall results demonstrate consistent English capabilities. This highlights the robustness of our method, which achieves cross-lingual knowledge editing while preserving the model’s original performance and significantly improving its proficiency in Chinese.
2 Analysis
2.1 Are both stages of X-KDE indispensable?
We examine the significance of the two stages in the X-KDE method through our ablation experiments, as shown in Table 4. {CJK*}UTF8gbsn Focusing solely on the improvement in performance metrics, Stage 1 undoubtedly plays a decisive role in our method, achieving significant gains (up to +25.64% average score) compared to the untrained baseline model. This stage enables cross-lingual knowledge editing via in-context learning, providing a substantial boost in model performance. While Stage 2 appears to offer a smaller improvement (a 2.25% average gain), a closer analysis highlights its practical importance. Stage 2 is particularly effective in adjusting the model’s output style to align with the target language, addressing issues such as incorrect language mixing (e.g., code-switching) or failure to generate responses in the expected linguistic format. For example, after Stage 1 updates the knowledge, the model may still produce a year like “2006” in the source language format. Stage 2 ensures the correct linguistic form, such as “2006年” in Chinese. The optimization of target-language preferences using ORPO not only improves factual accuracy but also ensures stylistic appropriateness in the target language. By refining the model’s preferences, ORPO helps it better adapt to the cultural and grammatical norms of the target language, addressing challenges like code-switching and maintaining consistency across multilingual contexts.
Methods Stages Score
Stage-1 Stage-2 en-avg zh-avg
Origin 62.41 64.96
X-KDE 88.05 86.99
77.14 66.38
91.04 88.49
Table 4: Ablation results on Bi-ZsRE benchmark with Llama2-7b-chat as the base model. The en-avg and zh-avg columns denote average scores when editing in English or Chinese, respectively.
2.2 Does every composition of the training data matter?
In this section, we focus on the composition of the training data. As shown in Table 5, the absence of any specific segment of the training data leads to a noticeable decline in editing performance, whether in monolingual or cross-lingual settings. Excluding either monolingual or cross-lingual training data causes a sharp drop in performance in the corresponding areas. When the in-scope data is omitted, the model tends to retain its original knowledge, resulting in reduced reliability, generality, and portability. On the other hand, removing the out-of-scope data causes the model to overly depend on the updated knowledge, thus diminishing locality. Similarly, removing the edit descriptors from the training data prevents the model from effectively utilizing the updated knowledge, leading to a drop in all metrics. Interestingly, training data with longer text samples seems to enhance the model’s comprehension and improve portability. In summary, each component of the training data plays a unique and indispensable role, and omitting any part negatively impacts the model’s performance across all key metrics.
Methods en-en en-zh
R. G. L. P. R. G. L. P.
X-KDE 99.93 99.87 90.15 76.41 94.81 94.65 95.05 77.43
-w/o mono. data 78.93 76.60 77.33 68.21 94.6 94.52 94.73 76.02
-w/o coss-lin. data 99.91 99.81 88.97 77.40 76.86 76.82 86.99 67.49
-w/o in-scope 81.02 82.58 93.93 69.15 75.17 74.85 93.56 69.62
-w/o out-of-scope 99.99 99.45 70.71 76.64 92.91 92.73 76.21 73.48
-w/o edit descriptor 87.53 81.99 67.69 66.53 84.26 84.12 79.43 74.13
-w/o long-text 100 99.82 93.54 73.63 92.35 92.75 93.16 72.46
Table 5: Ablation results in the monolingual and the cross-lingual setting, examining “reliability” (R.), “generality” (G.), “locality” (L.), and “portability” (P.).
2.3 Why choose ORPO as the preferred optimization method?
We evaluate several popular preference optimization methods using a held-out dataset from our training data, specifically direct policy optimization (DPO)rafailov2023direct, contrastive preference optimization (CPO)xu2024contrastive, Kahneman-Tversky Optimization (KTO)ethayarajh2024kto, and odds ratio preference optimization (ORPO)hong2403orpo on the Bi-ZsRE benchmark. As shown in Table 6, ORPO outperforms the other methods, achieving the best overall performance. While CPO and KTO also yield similar improvements, ORPO excels in preserving irrelevant target-language samples, demonstrating superior locality. In contrast, DPO results in a performance decline, which we attribute to the absence of negative log-likelihood (NLL) constraints, potentially weakening the model’s instruction-following capabilities. Based on these results, we adopt ORPO as the default optimization method for the second phase of our approach, as it provides the most significant improvement.
Method Eff. Gen. Loc.a Por. Avg.
SFT 90.22 90.2 89.22 64.03 83.41
     +DPO 88.47 88.3 89.18 61.11 83.26
     +CPO 92.41 92.67 90.97 67.09 85.78
     +KTO 92.23 92.01 89.22 67.23 85.17
     +ORPO 92.85 93.06 92.49 67.23 86.41
Table 6: Effects of different preference optimization methods with single edit setting on en-zh. Shades of cell color represent differences between preference optimization methods and simply SFT, where blue denotes better performance while red indicates worse.
Appendix A Details of Dataset Construction
A.1 Prompt Details of Sample Generation
Here, we present the detailed prompts for sample generation. Specifically, we guide LLMs in producing the queries and answers via the following prompts:
Query Generation Prompt [Edit description] will modify the knowledge. You must think according to the narrative of [Edit description].
[Edit description]: Who is Chris Klemmer affiliated with? University of Washington [Prompt]: Please generate a question related to Chris Klemmer. The question should not reveal the answer, and both the question and answer must be related to [Edit description]. [Generated Question]: Which university is Chris Klemmer associated with?
[Edit description]: What profession does Dagmar Lurz pursue? Film director
[Prompt]: Please generate a question related to Dagmar Lurz. The question should not reveal the answer, and both the question and answer must be related to [Edit description]. [Generated Question]: In which creative field has Dagmar Lurz displayed her extraordinary talent?
[Edit description]: When was Jana Fesslová born? April 20, 1977
[Prompt]: Please generate a question related to Jana Fesslová. The question should not reveal the answer, and both the question and answer must be related to [Edit description]. [Generated Question]: What major historical or cultural event in the 1970s aligns with Jana Fesslová’s birth date?
[Edit description]: What type of voice does Martin Crosby have? Contralto
[Prompt]: Please generate a question related to Martin Crosby. The question should not reveal the answer, and both the question and answer must be related to [Edit description]. [Generated Question]: What is Martin Crosby’s vocal type, and how does this type bring unique timbre and expressiveness to musical performances?
[Edit description]: Who directed "The Mentor"? Jean-Luc Godard
[Prompt]: Please generate a question related to "The Mentor". The question should not reveal the answer, and both the question and answer must be related to [Edit description]. [Generated Question]: Who directed "The Mentor," and how did they use their unique directorial style to shape the film’s artistic quality and influence?
[Edit description]: <instruct>
[Prompt]: Please generate a question related to <subject>. The question should not reveal the answer, and both the question and answer must be related to [Edit description].
[Generated Question]: <answer>
Answer Generation Prompt Imagine that the description of [Edit description] is correct. Please answer [question] based on [Edit description].
[Edit description]: Who is Chris Klemmer affiliated with? University of Washington [Question]: Who is Chris Klemmer affiliated with? [Prompt]: The answer should focus on Chris Klemmer, with University of Washington as a secondary focus, avoiding unnecessary details. [Answer]: Chris Klemmer is affiliated with the University of Washington. During his academic career at the University of Washington, he made significant achievements and contributions: 1.Academic Research: Klemmer conducted extensive research at the University of Washington, particularly in his field of expertise. His research findings not only advanced academic knowledge but also had a considerable impact on practical applications. 2.Teaching Contributions: As a professor at the University of Washington, Klemmer excelled in teaching and mentored many outstanding students. His teaching methods and course designs were well-received and inspired students to innovate. 3.Academic Leadership: Klemmer held various academic leadership positions at the University of Washington, such as department chair or research center director. In these roles, he promoted academic projects, fostered interdisciplinary collaboration, and enhanced the institution’s academic reputation. 4.Publications: Klemmer published numerous high-level academic papers and books, which received widespread attention and citation in the academic community, further solidifying his standing in academia.
[Edit description]: What profession does Dagmar Lurz pursue? Film director
[Question]: In which field has Dagmar Lurz shown exceptional talent? [Prompt]: The answer should focus on Dagmar Lurz, with film directing as secondary, avoiding unrelated content. [Answer]: Dagmar Lurz has demonstrated her exceptional talent in the field of film directing. As a film director, she is responsible for guiding actors’ performances and overseeing the film’s overall artistic style and narrative pacing. Her work involves interpreting scripts, arranging scenes, guiding cinematography, and making editing decisions, all requiring rich creativity and sharp artistic insight. Through her directorial work, Dagmar Lurz has left a unique mark on film art, offering viewers deep visual and emotional experiences.
[Edit description]: <instruct>
[Question]: <question>
[Prompt]: Please generate a question related to <subject>. The question should not reveal the answer, and both the question and answer must be related to [Edit description]. The answer should focus on <subject>, with <new answer> as secondary, avoiding unrelated content.
[Generated Answer]: <new answer>
Out-of-scope Generation Prompt In the following statements, "changed answer" represents the modified factual knowledge. When the answer is changed, other properties of the subject should remain unchanged. For example, if we edit basketball player Grant Hill to a football player, this would not affect his nationality. Therefore, for irrelevant attributes such as country, the output should remain consistent with the pre-edit version. You should recall an irrelevant attribute and generate a question and answer based on that irrelevant attribute and the "subject".
Question: Who is the father of Juan María Bordaberry?
Subject: Juan María Bordaberry Changed answer: Gabrielle Bordaberry Irrelevant attribute recalled: place of death New question: Where did Juan María Bordaberry die? New answer: Montevideo
Question: Who published the game Street Rod 2?
Subject: Street Rod 2 Changed answer: Sierra Entertainment Irrelevant attribute recalled: release format New question: What is the release format of Street Rod 2? New answer: Floppy disk
Question: What is the status of the Cross River Gorilla?
Subject: Cross River Gorilla Changed answer: Endangered Irrelevant attribute recalled: classification level New question: What is the classification level of the Cross River Gorilla? New answer: Subspecies"""
[Question]: <question>
[Subject]: <subject>
[Changed Answer]: <new answer>
A.2 Prompt Details of Quality Control
Moreover, we employ LLMs in Quality Controls. Judging and scoring via the following prompts:
Prompt for judging Please act as a fair judge and determine whether the [answer] answers the [question] based on the [Edit description]. Provide an explanation and strictly follow the format:
- If the answer is based on the edit description, output "[T]"
- If it is not, output "[F]".
[Edit description]: <instruct>
[Question]: <question>
[Answer]: <answer>
Prompt for scoring Please act as a fair judge and rate the sentence based on the following criteria:
1. Sentence complexity: Evaluate the complexity of the sentence, such as inversion, imperative sentences, sentences with word inflections, or sentences starting with multiple adverbs, nouns, and subjects. The more complex, the higher the score.
2. Vocabulary richness: Evaluate the diversity of vocabulary used. The more diverse, the higher the score.
3. Faithfulness: Evaluate whether the [answer] faithfully adheres to [Edit description], meaning it accurately answers [question]. If the answer highly matches the description, the score is higher; if the question leaks the answer, deduct points.
Provide separate scores for each criterion (1-10), and calculate the average score, then output the sentence with the highest score. The output format should be:
[Sentence complexity: score; Vocabulary richness: score; Faithfulness: score; Overall score: score]
[Edit description]: <instruct>
[Question]: <question>
[Answer]: <answer>
A.3 Training Data Statistics
Table 8 lists the statistics of our constructed high-quality dataset, which includes 388k samples and covers two languages: English (En) and Chinese (Zh). To prevent data leakage, we only sample instances from the training sets.
Appendix B Experimental Setup
B.1 Alignment Score Calculation
We employ NLLB-600M-distilled [^1] https://huggingface.co/facebook/nllb-200-distilled-600M as the multilingual translation model. And since the translation model is trained on a large-scale parallel dataset in both source and target languages ($X$ and $Y$) by maximizing the conditional generation probability $P(Y|X;\theta)$. Higher alignment between $Y$ and $X$ results in a higher conditional generation probability $P(Y|X)$. To compute the "alignment" score, we input the model’s response and force-decode the ground-truth $Y$ in target language.
B.2 Implementation Details
All experiments were executed on 8 NVIDIA A100 GPUs (80G). We employ EasyEdit wang2023easyedit to implement all the baselines with the default settings. We employ llamafactory zheng2024llamafactory to implement Cross-lingual Edition Instruction Tuning (XE-IT) phase of our method. When training on the English editing only subset, the duration is approximately 10 to 15 hours. Hyper-parameters of our X-KDE are in Table 7.
Hyperparameter XE-IT TL-PO
Learning rate 5e-6 1e-6
Max length 2560 1024
Optimizer AdamW AdamW
Scheduler cosine cosine
Weight decay 0.1 0.05
warmup steps 100 100
Table 7: Hyper-parameters for training our X-KDE.
Appendix C Used Scientific Artifacts
We list scientific artifacts used in our work blow. And we use of these existing artifacts is consistent with their intended use.
Data Source Lang. # in-scope # out-scope # Total Avg Token
w/ edit w/o edit w/ edit w/o edit
ZsRE En 20,000 20,000 20,000 20,000 80,000 48
Zh 20,000 20,000 20,000 20,000 80,000 84
HalluEditBench En 2,000 2,000 2,000 2,000 8,000 38
Zh 2,000 2,000 2,000 2,000 8,000 60
RIPPLEEDITS En 2,250 2,250 2,250 2,250 9,000 53
Zh 2,250 2,250 2,250 2,250 9,000 88
WikiBio En 250 250 250 250 1,000 162
Zh 250 250 250 250 1,000 294
MQUAKE En 4,000 4,000 4,000 4,000 16,000 266
Zh 4,000 4,000 4,000 4,000 16,000 334
COUNTERFACT En 7,500 7,500 7,500 7,500 30,000 530
Zh 7,500 7,500 7,500 7,500 30,000 888
Total En 36,000 36,000 36,000 36,000 144,000 170
Zh 36,000 36,000 36,000 36,000 144,000 266
Table 8: Statistics of our training data (Lang.:language). “Avg Token” denotes the average length(token-level) of samples, and “edit” indicates the edit descriptor.
Appendix D Supplemental Experiment Results
D.1 Detailed Results of Llama2-7b-chat
More detailed results on MzsRE of Llama2-7b-chat are listed in Table 9 and Table 10.
D.2 Detailed Results of Qwen2.5-7B-Instruct
To verify the universality of our method, we conducted experiments on Qwen2.5-7B-Instruct. More detailed results are listed in Table 11, Table 12 and Table 13.
Metrics Methods en-en en-cz en-de en-du en-es en-fr en-pt en-ru en-th en-tr en-vi en-zh en-avg
Reliability FT-L 52.92 41.81 39.79 39.02 39.49 39.72 39.26 39.79 36.44 36.86 46.21 51.81 41.93
FT-M 99.96 66.93 70.16 67.17 63.69 64.98 64.22 48.96 36.46 57.54 66.80 56.89 63.65
ROME 96.36 56.54 60.82 58.89 57.41 56.43 54.91 41.69 35.44 45.76 56.94 49.94 55.93
MEMIT 95.44 62.37 64.82 64.12 59.46 61.90 58.69 44.54 36.40 49.15 61.34 52.05 59.19
IKE 99.65 83.22 80.61 79.36 76.69 78.48 75.37 67.62 54.38 76.90 81.22 67.83 76.78
LTE 100.00 84.29 81.71 80.60 77.67 79.11 77.39 72.02 62.04 78.87 81.92 76.93 79.38
X-KDE 99.93 92.78 87.43 88.89 85.71 87.49 89.87 89.32 89.66 91.23 87.55 93.07 90.24
Generality FT-L 49.60 40.75 38.87 38.36 39.68 39.12 39.56 38.97 36.89 37.18 45.89 51.71 41.38
FT-M 95.53 65.45 68.15 65.09 62.39 62.28 61.63 47.69 36.88 56.87 65.97 56.52 62.04
ROME 85.13 54.99 58.91 56.99 56.58 54.47 53.94 40.68 35.36 45.06 56.38 50.31 54.07
MEMIT 89.59 60.71 63.80 61.98 58.10 59.40 57.63 43.31 36.77 48.68 60.51 52.01 57.71
IKE 99.54 82.67 80.78 79.18 76.37 78.22 75.49 67.51 54.26 76.97 80.99 67.88 76.65
LTE 99.87 84.26 81.63 81.07 77.51 78.99 77.38 71.46 61.90 78.26 81.37 76.24 79.16
X-KDE 99.68 92.87 87.25 88.87 85.16 87.57 89.93 89.10 89.21 91.25 87.62 93.11 90.14
Locality FT-L 93.96 90.78 81.06 88.98 83.32 89.30 90.98 89.53 90.18 88.95 93.02 85.56 88.80
FT-M 97.71 96.94 96.24 96.57 96.36 97.56 97.49 97.31 96.93 97.25 98.04 94.61 96.92
ROME 97.81 96.12 97.57 96.80 97.36 96.98 97.14 96.70 96.28 96.83 97.60 97.70 97.07
MEMIT 98.55 98.24 98.55 98.08 98.35 98.30 98.45 98.33 98.88 98.97 98.89 98.76 98.53
IKE 58.13 61.35 65.57 61.52 63.93 60.42 59.42 58.90 68.84 63.97 68.40 64.54 62.91
LTE 89.28 77.01 77.90 77.68 81.54 81.51 81.23 78.39 79.86 76.34 82.93 86.63 80.86
X-KDE 93.12 78.76 79.88 77.19 81.29 78.97 80.00 82.78 82.08 72.62 82.11 91.91 81.73
Portability FT-L 52.85 46.85 43.51 43.21 44.47 44.91 43.72 47.05 39.92 41.14 54.05 55.13 46.40
FT-M 57.17 48.66 46.38 46.34 47.09 47.54 46.36 48.41 38.55 42.50 55.53 52.16 48.06
ROME 58.46 49.79 48.58 47.06 48.29 48.83 47.30 49.21 38.11 42.38 56.62 51.81 48.87
MEMIT 57.02 50.41 47.96 47.26 47.26 48.47 47.21 49.25 38.56 43.44 57.16 52.19 48.85
IKE 70.97 56.44 58.87 56.91 58.05 58.41 56.33 57.96 39.46 48.69 62.86 58.97 56.99
LTE 77.29 61.85 64.91 63.82 61.53 62.83 62.39 61.43 44.51 51.04 65.37 67.47 62.04
X-KDE 76.13 60.53 58.74 56.94 55.19 58.85 58.81 62.89 56.18 48.69 61.78 74.04 60.73
Table 9: Results on MzsRE dataset for editing performed in English using Llama2-7b-chat. Here, “en-zh” means that English serves as the source language and Chinese as the target language, with similar interpretations for the other pairs. “en-avg” denotes the average performance across cross-lingual scenarios.
Metrics Methods zh-en zh-cz zh-de zh-du zh-es zh-fr zh-pt zh-ru zh-th zh-tr zh-vi zh-zh zh-avg
Reliability FT-L 40.81 38.16 36.21 35.60 36.28 36.45 35.55 38.88 33.98 34.50 43.32 54.79 38.71
FT-M 51.87 48.51 46.71 45.70 45.69 45.98 45.65 46.97 39.44 44.74 54.06 100.00 51.28
ROME 44.15 40.06 38.04 37.62 38.44 37.99 37.69 39.25 32.94 36.49 44.82 73.48 41.75
MEMIT 51.87 41.45 39.61 39.29 39.19 39.23 38.78 40.85 33.77 38.49 46.72 76.12 43.78
IKE 65.88 68.68 67.63 66.75 65.06 65.63 63.82 63.52 52.39 61.32 70.90 99.85 67.62
LTE 65.44 64.74 62.05 62.91 61.09 60.85 61.20 63.09 55.71 58.15 67.02 99.76 65.17
X-KDE 94.64 84.40 83.05 81.08 80.33 81.22 83.38 82.56 83.09 78.69 81.47 99.99 84.49
Generality FT-L 40.67 37.70 36.35 35.18 36.60 35.49 35.67 38.19 33.86 34.97 43.14 53.90 38.48
FT-M 51.24 48.24 46.49 45.30 45.71 45.63 45.81 46.32 39.62 45.06 54.42 99.68 51.13
ROME 43.80 39.72 38.01 37.83 38.26 36.74 38.62 38.46 32.76 36.46 45.02 71.13 41.40
MEMIT 51.24 41.16 40.14 38.22 39.10 38.80 39.06 40.15 34.18 38.22 46.34 74.21 43.40
IKE 65.75 67.90 67.48 66.39 65.01 65.35 63.50 63.44 52.72 61.03 70.27 99.28 67.34
LTE 64.94 64.53 62.72 62.31 61.15 60.50 61.11 62.94 55.39 58.29 66.88 98.69 64.95
X-KDE 94.51 84.27 82.43 81.46 80.12 81.08 82.69 82.19 82.81 78.33 81.07 98.89 84.15
Locality FT-L 94.81 89.42 83.41 88.81 83.09 89.92 90.09 86.63 79.94 85.75 89.69 66.38 85.66
FT-M 98.19 96.70 92.82 95.79 93.90 97.68 97.17 96.46 93.05 95.49 97.18 79.74 94.51
ROME 97.93 96.17 97.41 96.23 96.81 96.66 96.99 95.89 94.36 96.15 97.23 96.42 96.52
MEMIT 98.19 98.05 98.52 98.68 98.76 98.52 98.50 97.51 96.55 98.10 98.36 96.06 97.98
IKE 69.41 63.74 63.22 64.02 62.24 61.69 62.10 61.15 67.15 64.73 69.31 67.91 64.72
LTE 89.26 76.59 80.09 78.01 81.86 81.44 81.22 80.31 80.24 76.63 82.92 86.67 81.27
X-KDE 94.07 80.53 81.79 78.13 83.49 81.18 81.72 84.22 84.14 75.66 83.15 92.20 83.36
Portability FT-L 55.25 47.54 45.76 44.73 46.21 46.61 44.96 48.20 37.99 41.16 54.13 48.07 46.72
FT-M 55.30 48.13 45.78 45.25 46.09 47.04 45.22 48.75 40.32 42.12 55.57 61.79 48.45
ROME 52.66 47.19 44.60 44.89 44.56 45.64 44.11 47.96 36.40 40.23 53.87 48.34 45.87
MEMIT 55.30 47.78 45.99 45.60 46.00 46.91 45.32 48.10 37.82 41.51 54.88 50.83 47.17
IKE 63.06 54.77 58.40 56.25 55.59 56.72 54.58 57.65 40.80 47.15 61.58 66.44 56.08
LTE 68.30 59.86 60.92 59.62 58.27 59.71 58.59 60.03 45.83 48.53 63.11 69.40 59.35
X-KDE 67.95 60.34 59.10 58.10 55.89 58.43 59.37 61.73 56.27 49.69 62.78 73.43 60.26
Table 10: Results on MzsRE dataset for editing performed in Chinese using Llama2-7b-chat. Here, “zh-en” means that Chinese serves as the source language and English as the target, with similar interpretations for the other pairs. “zh-avg” denotes the average performance across cross-lingual scenarios.
Method Test in English Test in Chinese
Reliability Generality Locality Portability Reliability Generality Locality Portability Avg.
Edit in English
FT-L 62.89 65.93 75.06 38.83 44.25 44.36 68.35 40.21 54.98
FT-M 100.0 99.35 92.44 52.59 64.03 63.61 88.48 53.22 76.72
ROME 99.48 92.83 98.63 56.69 58.91 58.44 98.47 54.93 77.30
MEMIT 96.77 89.30 98.78 55.60 59.97 59.35 98.49 54.31 76.57
IKE 97.32 98.56 50.26 67.30 69.89 69.45 57.07 57.94 70.97
LTE 99.78 99.28 87.64 74.23 73.95 74.40 84.34 61.85 81.93
X-KDE(Ours) 99.72 99.56 88.79 73.96 90.42 90.20 91.53 62.59 87.10
Edit in Chinese
FT-L 32.59 33.17 73.39 37.89 48.85 53.49 54.88 28.17 45.30
FT-M 53.31 52.80 92.53 51.52 100.0 99.85 79.38 52.84 72.78
ROME 45.66 45.31 98.31 52.74 99.36 94.77 97.97 57.06 73.90
MEMIT 45.68 44.25 98.94 52.26 98.07 94.20 96.69 57.70 73.47
IKE 79.59 78.77 49.57 65.20 96.37 96.47 66.05 61.59 74.20
LTE 79.40 78.50 86.64 70.24 98.95 98.60 84.54 64.53 82.68
X-KDE(Ours) 94.78 94.77 95.14 67.50 99.79 98.29 90.57 61.30 87.77
Table 11: Cross-lingual editing performance of different methods on Qwen2.5-7B-Instruct backbones. Results in green indicates the best results. “Avg.” represents the overall mean of all metrics evaluated across the two languages.
Metrics Methods en-en en-cz en-de en-du en-es en-fr en-pt en-ru en-th en-tr en-vi en-zh en-avg
Reliability FT-L 63.77 50.88 50.30 47.23 47.40 51.08 48.18 42.59 44.37 47.86 48.39 44.91 48.91
FT-M 100.0 71.56 74.67 70.47 66.45 68.94 68.82 57.35 51.57 67.36 64.79 64.59 68.88
ROME 99.44 55.82 63.27 61.38 57.41 59.44 60.23 49.78 48.74 53.73 53.06 59.23 60.13
MEMIT 96.92 54.87 61.36 58.79 54.67 58.15 58.02 49.53 47.71 52.49 51.88 60.27 58.72
IKE 97.89 82.71 82.84 78.05 76.34 79.26 78.69 69.95 66.43 77.28 75.93 70.53 77.99
LTE 99.70 84.28 84.73 80.76 78.02 82.07 80.40 76.52 69.53 81.06 77.89 73.35 80.69
X-KDE 98.56 89.60 86.69 86.94 84.84 84.87 85.30 89.05 87.61 90.00 79.47 89.74 87.72
Generality FT-L 66.81 50.62 50.42 46.74 47.63 51.18 48.35 42.85 44.81 47.66 47.50 44.91 49.12
FT-M 99.26 70.46 73.67 69.37 65.68 67.06 66.14 56.27 51.83 65.35 62.55 64.09 67.64
ROME 93.67 54.71 61.05 58.82 55.48 57.48 58.01 48.59 48.32 51.49 51.57 59.13 58.19
MEMIT 90.32 54.69 59.12 56.52 53.93 55.96 55.53 48.25 47.77 51.16 50.64 59.72 56.97
IKE 98.52 82.75 82.71 77.83 75.86 78.92 78.30 69.57 66.83 77.26 75.26 70.68 77.87
LTE 99.30 84.48 84.57 80.39 77.93 81.49 80.34 76.56 69.45 81.07 77.56 73.63 80.56
X-KDE 98.00 89.91 86.55 87.07 84.88 84.53 85.49 89.03 87.49 89.82 79.42 89.95 87.68
Locality FT-L 74.68 75.09 62.23 69.32 62.46 70.98 73.56 74.63 79.92 66.58 75.33 67.88 71.06
FT-M 92.46 90.64 83.27 87.98 83.72 90.23 90.39 90.89 93.08 86.20 91.12 88.30 89.02
ROME 98.71 97.43 97.32 98.05 98.35 97.37 98.41 98.49 98.46 97.53 98.53 98.35 98.08
MEMIT 98.76 98.49 98.21 98.49 98.77 98.63 98.48 98.74 98.75 98.59 98.47 98.51 98.57
IKE 50.52 55.71 57.39 53.66 57.51 58.40 56.95 61.17 65.64 59.46 60.03 57.45 57.82
LTE 88.28 80.06 81.15 78.18 82.89 84.62 82.80 86.32 85.12 78.61 80.00 84.74 82.73
X-KDE 95.22 77.82 87.16 77.71 79.86 83.98 82.35 87.42 86.40 75.54 80.53 92.59 83.88
Portability FT-L 38.08 38.77 36.92 36.49 37.42 39.98 40.22 45.27 47.89 39.07 35.92 40.23 39.69
FT-M 52.17 49.42 48.36 46.27 46.92 50.18 49.86 53.84 54.22 46.93 45.66 52.83 49.72
ROME 56.40 49.68 49.33 48.00 49.17 53.04 51.94 54.93 54.53 47.56 46.14 54.95 51.31
MEMIT 54.82 49.69 49.62 47.80 49.48 51.74 51.67 55.01 54.57 48.15 46.94 54.32 51.15
IKE 67.00 55.75 59.25 55.11 56.31 60.00 58.76 59.04 55.32 52.44 52.16 58.27 57.45
LTE 74.25 61.72 65.09 61.27 62.63 65.92 64.21 65.30 60.78 59.96 57.94 61.36 63.37
X-KDE 70.36 53.71 54.93 52.23 54.76 56.20 56.57 60.27 58.51 52.41 49.23 61.18 56.70
Table 12: Results on MzsRE dataset for editing performed in English using Qwen2.5-7B-Instruct. Here, “en-zh” means that English serves as the source language and Chinese as the target language, with similar interpretations for the other pairs. “en-avg” denotes the average performance across cross-lingual scenarios.
Metrics Methods zh-en zh-cz zh-de zh-du zh-es zh-fr zh-pt zh-ru zh-th zh-tr zh-vi zh-zh zh-avg
Reliability FT-L 33.24 34.37 32.47 31.74 30.58 33.23 32.99 35.73 40.39 29.41 28.90 48.62 34.31
FT-M 52.97 51.26 51.88 49.22 48.13 49.82 51.21 52.25 51.33 50.00 47.02 100.00 54.59
ROME 46.29 42.41 43.29 42.32 41.15 42.64 43.13 45.94 47.14 44.22 40.61 99.51 48.22
MEMIT 46.36 42.81 44.02 42.29 40.87 42.69 43.75 46.06 46.78 43.18 41.53 98.59 48.25
IKE 80.22 71.25 76.32 69.69 69.28 71.10 70.64 66.51 65.03 68.96 66.66 96.32 72.67
LTE 79.54 71.83 75.40 71.18 67.28 70.02 68.88 72.39 66.11 70.15 68.79 99.03 73.38
X-KDE 94.97 79.86 84.24 79.12 77.97 81.24 79.16 75.77 67.75 77.76 76.48 99.85 81.18
Generality FT-L 32.90 34.32 32.80 31.56 31.00 33.88 33.40 35.69 40.79 29.89 29.57 53.08 34.91
FT-M 52.66 51.13 52.15 49.12 47.66 48.95 50.46 51.47 51.51 50.22 46.62 99.96 54.33
ROME 46.11 42.31 42.60 41.57 40.65 41.88 42.77 44.60 46.34 42.97 40.80 95.23 47.32
MEMIT 45.03 42.58 42.79 41.50 40.83 41.81 42.51 45.20 46.48 42.65 41.33 95.06 47.31
IKE 78.68 72.05 76.01 69.73 69.72 71.07 71.09 66.08 65.18 69.40 66.86 96.50 72.70
LTE 78.20 71.15 75.18 70.72 66.41 69.68 69.22 72.32 66.04 69.93 68.50 98.34 72.97
X-KDE 94.93 79.44 84.67 79.08 77.85 81.37 79.20 75.49 68.11 77.53 76.29 98.51 81.04
Locality FT-L 73.25 67.27 58.50 63.17 58.05 65.53 68.75 67.52 69.27 56.25 64.19 54.75 63.87
FT-M 92.53 89.49 83.74 87.61 85.73 89.50 90.45 88.96 90.36 83.69 90.04 79.06 87.60
ROME 98.26 97.35 97.07 96.83 97.80 97.37 97.90 97.94 97.92 96.74 97.54 97.90 97.55
MEMIT 98.94 98.36 98.42 97.94 98.49 98.69 98.81 98.36 97.69 97.96 97.76 96.67 98.18
IKE 50.43 57.50 57.28 53.59 55.05 56.53 56.76 61.92 68.25 61.89 61.34 65.49 58.84
LTE 87.33 79.75 81.18 76.95 83.05 83.28 80.42 85.36 84.31 78.27 80.59 84.56 82.09
X-KDE 93.87 81.90 84.70 81.11 86.56 87.89 87.14 88.08 90.18 79.84 84.63 90.42 86.36
Portability FT-L 37.03 36.69 34.82 34.70 34.97 37.81 37.40 41.47 44.90 33.05 30.24 27.71 35.90
FT-M 50.78 47.32 45.81 44.66 45.38 47.77 47.23 51.59 53.59 44.78 42.89 51.91 47.81
ROME 51.70 47.48 47.36 45.42 46.54 50.09 49.16 54.37 54.39 46.86 44.10 56.78 49.52
MEMIT 51.54 47.74 47.22 45.78 46.19 49.21 48.87 54.48 54.24 46.25 44.32 57.65 49.46
IKE 64.32 56.14 58.98 55.44 56.10 59.75 58.63 61.12 57.28 53.05 52.57 61.75 57.93
LTE 70.11 61.02 64.45 60.44 62.24 65.28 63.82 65.28 60.67 59.51 57.18 64.36 62.86
X-KDE 67.44 56.98 61.07 58.20 60.59 61.93 61.62 63.60 58.95 55.87 55.28 60.92 60.20
Table 13: Results on MzsRE dataset for editing performed in Chinese using Qwen2.5-7B-Instruct. Here, “zh-en” means that Chinese serves as the source language and English as the target, with similar interpretations for the other pairs. “zh-avg” denotes the average performance across cross-lingual scenarios.